OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 4/5, what is the equation of the line?

OpenStudy (anonymous):

The equation can be the y intercept equation of a line y = mx + b b = y-intercept = (0, 1) b = 1 m = slope can you show me the equation?

OpenStudy (anonymous):

\[\frac{ y-y _{1} }{x-x _{1} } = m\] \[\frac{ y-1 }{ x-0 } = \frac{ 4 }{ 5 }\] \[5y-5=4x\] so equation of line will be \[4x-5y+5=0\]

OpenStudy (butterflydreamer):